OpenStudy (anonymous):

2sin^2u+sinu-6=0 What is the trick on factoring this?

OpenStudy (anonymous):

factor

OpenStudy (anonymous):

yeah I got that, but how?

OpenStudy (anonymous):

omg, im really having a brain fart today...

OpenStudy (anonymous):

2x^2+x-6 = (2x-3)(x+2) (2sinu-3)(sinu+2) = 0 sinu =-2 or sinu = 3/2

OpenStudy (anonymous):

thank you